2 - 5 For each system below, (a) solve the initial value problem, and (b) determine the type and stability of the critical pox= 5x - 5x2 X2 = 2xı + 3x2 xi(–117) = 7, x2(–117) = 3.
Answer #1

Civen that x;=52,522 x = 20, +34 x,(-117) 37 22(117) = 3 al Rewrite xi - 5.31 - 5*2 = 0 X-322 -27,50 (0-52*4- 5 X2 50 - (0-3)-117. Cute . Now x (+117) = 7 7 = {[%7C1*57c) ***84? C. (1-51) 31 (4837 4 226-117)= 3 - 328i el Canto - tech.s) Solving buthI hope it will be helpful for you if any query comment please if satisfied rating please